9). A small bag of sand is released from an ascending hot‑air balloon whose constant, upward velocity is ?0=2.75 m/s. Knowing that at the time of the release the balloon was 58.3 m above the ground, determine the time ? it takes for the bag to reach the ground from the moment of its release. Use  ?=9.81 m/s2.
r=

10)

You throw a baseball directly upward at time ?=0 at an initial speed of 14.1 m/s. What is the maximum height the ball reaches above where it leaves your hand? Ignore air resistance and take ?=9.80 m/s2.

maximum height: 10.143

At what times does the ball pass through half the maximum height?

earlier time at half maximum height:——
later time at half maximum height——-

11)

While standing at the edge of the roof of a building, a man throws a stone upward with an initial speed of 5.71 m/s. The stone subsequently falls to the ground, which is 11.3 m below the point where the stone leaves his hand.

At what speed does the stone impact the ground? Ignore air resistance and use ?=9.81 m/s2 for the acceleration due to gravity.

impact speed:——-

How much time is the stone in the air?

elapsed time:———

m

At what times does the ball pass through half the maximum height?

earlier time at half maximum height:
Answer #1

tvel Qq hat downward speed is upward speed is So, % = -2- 75 m/s to ocoles & & avel S = 58-3 mi a 9.81 m/s2 tog 102 = 8 + 2as